# Numerical analysis for leaky-integrate-fire networks under Euler–Maruyama

### 11 Mar 2026 v1

@@ 8 Discussion
From a modeling perspective, the transversality variable $A=I-I_{\mathrm{th}}$ has a direct interpretation. Large $A$ corresponds to robust threshold crossings, while small $A$ corresponds to fluctuation-driven, nearly tangential spikes that are intrinsically harder to resolve on a grid. The logarithmic inverse-moment bound shows that, under a boundary-flux control, these difficult crossings are rare enough that the good-set mean-square error remains $h$ up to polylogarithmic factors. In other words, the classical Euler–Maruyama $1/2$ strong behavior survives in the bulk, but resets leave a measurable numerical signature through threshold geometry. This is especially relevant for cortical regimes in which irregular spiking and weakly driven threshold crossings are part of the modeled phenomenon rather than a nuisance [12, 6].
@@
The main mathematical limitation of the current strong theory is that it treats $T$ and $L$ as fixed and uses a deliberately conservative recursion. Once an upstream spike-time error appears, the bound passes its maximal downstream impact forward through all later layers. This is robust for finite $T$ and $L$, but it ignores spike causality. A perturbation in layer $1$ changes the current in layer $2$ immediately, yet it can alter spike times in deep layers only through a cascade of later threshold crossings. Thus, for large depth or long time, the relevant object is not an instantaneous all-to-all transfer of the upstream supremum error, but a causal genealogy of descendant spikes. A sharper strong-error theory should therefore replace the present layerwise max recursion by a causal Green’s-function or Volterra representation for the hybrid variational equation. One natural target would be an estimate of the form

### 12 Mar 2026 v2

@@ 8 Discussion
From a modeling perspective, the transversality variable $A=I-I_{\mathrm{th}}$ has a direct interpretation. Large AA corresponds to robust threshold crossings, while small AA corresponds to fluctuation-driven, nearly tangential spikes that are intrinsically harder to resolve on a grid. The logarithmic inverse-moment bound shows that, under a boundary-flux control, these difficult crossings are rare enough that the good-set mean-square error remains $h$ up to polylogarithmic factors. In other words, the classical Euler–Maruyama $1/2$ strong behavior survives in the bulk, but resets leave a measurable numerical signature through threshold geometry. This is especially relevant for cortical regimes in which irregular spiking and weakly driven threshold crossings are part of the modeled phenomenon rather than a nuisance [12, 6].
@@
The main mathematical limitation of the current strong theory is that it treats $T$ and $L$ as fixed and uses a deliberately conservative recursion. Once an upstream spike-time error appears, the bound passes its maximal downstream impact forward through all later layers. This is robust for finite $T$ and $L$, but it ignores spike causality. A perturbation in layer 11 changes the current in layer 22 immediately, yet it can alter spike times in deep layers only through a cascade of later threshold crossings. Thus, for large depth or long time, the relevant object is not an instantaneous all-to-all transfer of the upstream supremum error, but a causal genealogy of descendant spikes. A sharper strong-error theory should therefore replace the present layerwise max recursion by a causal Green’s-function or Volterra representation for the hybrid variational equation. One natural target would be an estimate of the form